The integral factor of (x²y)dx- (x³ + y³)dy = 0 is


A) -y^4


B) -1/y^4


C) y^4 D) 1/y^4


1
Expert's answer
2021-08-23T04:48:30-0400

Using homogeneous method

"let \space y=vx\\

⇒\frac{dy}{dx}=\frac{(x³+y³)}{x²y}\\

\frac{dy}{dx}=\frac{(x³+ (vx)³)}{x²(vx)}\\

⇒ \frac{dy}{dx}=\frac{(x³+v³x³)}{x³v}\\

\frac{dy}{dx}=\frac{x³(1+v³)}{x³v}\\

\frac{dy}{dx}=\frac{(1+v³)}{v}\\

There for \frac{dy}{dx}=\frac{1+vdv}{dx}\\

⇒ \frac{1+vdv}{dx}=\frac{(1+v³)}{v}

⇒\frac{vdv}{dx}=\frac{(1+v³)}{v(-1)}\\

\frac{dy}{dx}=(\frac{1}{v+v²})-1\\

\frac{dy}{dx}=(1/v²+v-1/v)\\

⇒\frac{dv}{(v²+v∧-1-v)}=dx\\

⇒ (\frac{v³}{3})+(\ln v)-(\frac{v²}{2})=x\\

but v=\frac{y}{x}\\

⇒ \frac{(\frac{y}{x})³}{3} + (\ln(\frac{y}{x})) - \frac{(\frac{y}{x})²}{2}=x+c"